The current I varies directly as the electromotive force E and inversely as the resistance R. If a
current of 33 amperes flows through a system with 22 ohms resistance and electromotive force of
220 volts, what is the equation of variation? Find the current if the electromotive force of 200 volts with
the same resistor will send through the system.

Answer by Alan3354(69443)

The current I varies directly as the electromotive force E and inversely as the resistance R. If a current of 33 amperes flows through a system with 22 ohms resistance and electromotive force of 220 volts, what is the equation of variation? Find the current if the electromotive force of 200 volts with
the same resistor will send through the system.
[sic]
============================
E = I*R
220 volts will not give a current of 33 amps thru a 22 ohm resistance.
33*22 = 726 volts
----------------
E = I*R
I = E/R
I = 200/33 amperes